OpenStudy (anonymous):

3/21 divide by 6/7 help plz

TheSmartOne (thesmartone):

\[\frac{ 3 }{ 21 }\div \frac{ 6 }{7 }=\frac{ 3 }{ 21 }\times \frac{ 7 }{6 }=\frac{ 1 }{ 6 }\]

TheSmartOne (thesmartone):

@edkasugifdfhdfhd

OpenStudy (anonymous):

thank you so much

TheSmartOne (thesmartone):

np. Here look at these because it seems you are new here. One of them tells you what you can do on OS. And the other is for choosing one of them as a profile pic. Of course you don't have to take one of the owls for a profile pic. http://openstudy.com/study#/updates/543f2836e4b0ab604fb452de http://openstudy.com/study#/updates/5445acd8e4b051a66734d35b

TheSmartOne (thesmartone):

@edkasugifdfhdfhd